OpenStudy (anonymous):

A guitar has a power output of 0.25 W. If the sound waves emitted by the guitar travel in spherical wavefronts, what will be the intensity of the sound at a distance of 4 m?

OpenStudy (anonymous):

http://hyperphysics.phy-astr.gsu.edu/hbase/acoustic/invsqs.html That should give you a nicer explaination than I'd be able to write out on the spot, but feel free to ask if you have any other questions!

OpenStudy (anonymous):

umm please write @LifeEngineer

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

okay ill show you my work sorry for being lazy

OpenStudy (anonymous):

4*pi*2^2=50.26 0.25/50.26=4.9*10^-3

OpenStudy (anonymous):

@LifeEngineer

OpenStudy (anonymous):

Close! The distance from the source is 4m, not 2m, so you would end up with: \[I = \frac{ 0.25 }{ 4\pi 4^2 }\]

OpenStudy (anonymous):

That will give you an answer in watts per square meter. If you need an answer in decibels, then you can convert fairly easily using \[I(dB)=10\log( \frac{I}{I_o})\]

OpenStudy (anonymous):

hint: if you scroll down on the link I gave you, there is also a calculator for this type of problem, after it gives the explanation of how everything works